A Rocket has a mass given by m-ct where it has hurt for a time t, the relative backward velocity of the gass is Vr ol if it burns out a mass 1/2 where it is fired vertically from rest neglecting air resistance find the height reached and velocity when it is burn out completely.


1
Expert's answer
2021-11-04T10:17:25-0400

"v=v_r\frac{me^{-ct}}{0.5m}=2v_re^{-ct},"

"h=\frac{v^2}{2g}=\frac{2v_r^2e^{-2ct}}g."